MRs. B gave her students a clue to her age. In 12 years, I will be 3 times the age I was 32 years ago. How old is Mrs. B?
Answer by richwmiller(17219)

x+12=3*(x-32)
x+12=3x-96
108=2x
x=54
check
66=3*(22)
ok